D Question 4 1 pts Use the graphs of the two parabolas to answer the following question. Solve the system y- ar+bx+e (S, T P. R) Choose the correct answer
Answer #1

The given parabolas i.e. y = ax2+bx+c and y = dx2+ex+f intersect at the points (S,T) and (P,R). Further, the points of intersection are the solutions to this system of quadratic equations. Hence the 3rd option, i.e. (S,T) (P,R) is the correct answer..